what is the final temperature of a 50.00 g sample of ice at -10.00 C if it is exposed to 59.75 J of energy and remains ice?
1
Expert's answer
2017-02-06T07:20:14-0500
Isobaric mass heat capacity of ice is c = 2.05 J*g-1*K-1.
Q = c*m*(T2-T1)
T2 = Q/(c*m) + T1 = 59.75/(2.05*50.00) – 10.00 = -9.42 oC